OpenStudy (anonymous):

Find the magnitude & direction of the vector JK given points J(1,0) & K(-2,3).

OpenStudy (lgbasallote):

\[\huge \text{magnitude} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\huge \text{direction} = \tan^{-1} \left(\frac{y_2 - y_1}{x_2 - x_1} \right)\]